Find the inverse of matrix A given below using the formula A -1 =CT/│A│

A="\begin{vmatrix}
 -6 & -4 & -4 \\
 -1 & -4 & 6 \\\
-6 & 5 & -1
\end{vmatrix}"


1
Expert's answer
2022-03-18T09:35:37-0400

"A=\begin{vmatrix}
 -6 & -4 & -4 \\
 -1 & -4 & 6 \\
 -6 & 5 & -1
\end{vmatrix}"

Main determinant

"|A|=-6*((-4)*(-1) - 5*6) - (-1)*((-4)*(-1) - 5*(-4)) + -6*((-4)*6 - (-4)*(-4)) = 420"

The determinant is non-zero, therefore, the matrix is non-degenerate and it is possible to find the inverse matrix A-1 for it.

The inverse matrix will look like this:

"A^{-1}=\frac{1}{420}\begin{pmatrix}
 A_{11} & A_{12} & A_{13} \\
 A_{21} & A_{22} & A_{23} \\
 A_{31} & A_{32} & A_{33} \\
\end{pmatrix}"

where Aij are algebraic additions.

The transposed matrix.

"A_{11}=((-4)*(-1) - 6*5) = -26\\
A_{12}= -((-4)*(-1) - (-4)*5) = -24\\
A_{13}= ((-4)*6 - (-4)*(-4)) = -40\\
A_{21}= -((-1)*(-1) - 6*(-6)) = -37\\
A_{22}= ((-6)*(-1) - (-4)*(-6)) = -18\\
A_{32}= -((-6)*6 - (-4)*(-1)) = 40\\
A_{31}= ((-1)*5 - (-4)*(-6)) = -29\\
A_{32}= -((-6)*5 - (-4)*(-6)) = 54\\
A_{33}= ((-6)*(-4) - (-4)*(-1)) = 20\\"

"A^{-1}=\frac{1}{420}\begin{pmatrix}
 -26 & -24 & -40 \\
 -37 & -18 & 40 \\
 -29 & 54 & 20 \\
\end{pmatrix}"

"A^{-1}=\begin{pmatrix}
 \frac{-13}{210} & \frac{-2}{35} & \frac{-2}{21} \\
 \frac{-37}{420} & \frac{-3}{70} & \frac{2}{21} \\
 \frac{-29}{420} & \frac{9}{70} & \frac{1}{21} \\
\end{pmatrix}"